OpenStudy (ilovecake):

A recipe that makes 3 servings calls for 3/5 cup of apple juice and 3/4 cup of cranberry juice. How many cups of juice are there per serving? A. 4 1/20 B. 1 7/20 C. 2/3 D. 9/20

OpenStudy (ilovecake):

Medal and fan

OpenStudy (ilovecake):

@AccessDenied

OpenStudy (ilovecake):

someone help

OpenStudy (anonymous):

is this a test quiz or homewok

OpenStudy (ilovecake):

Homework @P!nkfanlol444

OpenStudy (anonymous):

OK I can help but not give answers

OpenStudy (ilovecake):

Ok

OpenStudy (anonymous):

so what do you know from the question

OpenStudy (ilovecake):

A recipe that makes 3 servings calls for 3/5 cup of apple juice and 3/4 cup of cranberry juice.

OpenStudy (ilovecake):

That is what I know

OpenStudy (anonymous):

ok

OpenStudy (ilovecake):

Do I add the numbers?

OpenStudy (anonymous):

we need first find a common denominator..a number both 5 and 4 go into evenly..20 seems about right so we need to multiply the whole first fraction by 4 to make the denominator 20...and the whole second fraction by 5 to make that denominator 20

OpenStudy (ilovecake):

So C is not one of the right answers

OpenStudy (anonymous):

4/4x3/5+5/5x3/4 this will come out 12/20+15/20 whcih will equal............. Hint add 15/20 + 12/20 and it will give you your answer

OpenStudy (ilovecake):

27

OpenStudy (anonymous):

3/5 + 3/4 = 12/20 + 15/20 = 27/20 (total juice) 3/(27/20) = 1/ (x) --- 3 servings to 27/20 juice = 1 serving to x juice cross multiply (3)(x) = (27/20)(1) 3x = 27/20 -- divide by 3 x = (27/20) / 3 x = 27/20 * 1/3 x = 27/60 which reduces to 9/20

OpenStudy (anonymous):

so your answer is 9/20 ;)

OpenStudy (ilovecake):

Ok I understand it now thanks so much :)

OpenStudy (anonymous):

yup welcom to openstudy @llovercake

OpenStudy (ilovecake):

Lovercake lol

OpenStudy (anonymous):

ohh lol